OpenStudy (anonymous):

Factor. x2 + 6x + 9 (Points : 1) (x + 6)^2 (x – 6)^2 (x + 3)^2 (x – 3)^2

OpenStudy (whpalmer4):

The answer for the previous problem is incorrect, we should go back and work that first!

OpenStudy (anonymous):

alright so whats the answer for this one @whpalmer4

OpenStudy (whpalmer4):

well, let's go through it, same process: \[x^2+6x+9\]\[ax^2+bx+c\]\[a=1,b=6,c=9\] What is \(a*c=\)?

OpenStudy (anonymous):

9x^2.?

OpenStudy (whpalmer4):

no, we are just multiplying the coefficients, which are the numbers in front of the variables. as I said, \[a = 1, b = 6, c = 9\]\[a*c = \]

OpenStudy (anonymous):

9

OpenStudy (whpalmer4):

Very good. Now what are the two factors of 9 that add up to 6?

OpenStudy (whpalmer4):

here are all the factors of 9: 1*9 = 9 3*3 = 9 -1*-9 = 9 -3*-3 = 9 which pair of factors adds up to 6?

OpenStudy (whpalmer4):

@dsfasdfasdfga are you going to finish this problem, or not?